OpenStudy (anonymous):

how do i factor x^3 -3x -2 = 0? which method do i use?

sam (.sam.):

are you allowed to use calculator for test?

OpenStudy (anonymous):

its not a test..i can use anything but i'd rather do it by hand...i need to know how to work it..

OpenStudy (anonymous):

i just need the next step..just a hint please

sam (.sam.):

you need to find a zero of the polynomial f(x)=x^3 -3x -2 The fastest way is to use trial an error Try x=2, f(2)=2^3-3(2)-2 f(2)=0 Then, (x-2) is a factor of f(x) Then do long division (look at the attachment) you'll get x^2+2x+1 (x^2+2x+1)(x-2) Factor x^2+2x+1 (x+1)^2(x-2)

OpenStudy (anonymous):

that 0 that's taking the x^2 spot..is that supposed to be a one?

OpenStudy (anonymous):

^ doubt it when factoring normally you set it to zero...

OpenStudy (anonymous):

get a, b, c terms on one side to get ready to factor. it leaves the right side a 0 if done right.

OpenStudy (anonymous):

actually wait what? restate your question?

OpenStudy (anonymous):

i've got it now..thank you guys..

OpenStudy (anonymous):

Great!